OpenStudy (anonymous):

Will fan and medal!! HELP!!! Sally initially has 4 hours of pop music and 3 hours of classical music in her collection. Every month onwards, the hours of pop music in her collection is 8% more than what she had the previous month. Her classical music does not change. Which function shows the total hours of music she will have in her collection after x months? f(x) = 4(1.08)x + 3 f(x) = 3(1.08)x + 4 f(x) = 4(0.08)x + 3 f(x) = 3(0.08)x + 4

OpenStudy (anonymous):

i think it is C, but I'm not sure @HELP!!!! @acxbox22

OpenStudy (acxbox22):

it is A because the 3 represents the classical and is not changing and the pop gows by 8%, so 1 +0.08=1.08 which leaves us with the equation y=4(1.08)^x+3

OpenStudy (anonymous):

Took the words right out of my mouth! It's A

OpenStudy (anonymous):

Okay thank you soo much!!! but one more question

OpenStudy (anonymous):

okay :)

OpenStudy (anonymous):

fifth root of 8 multiplied by square root of 8 over cube root of 8 to the power of 5

OpenStudy (anonymous):

\[\frac{ 5^{\sqrt{8}*\sqrt{8}} }{ 3\sqrt{8^5} }\]

OpenStudy (anonymous):

Is that a multiple choice question?

OpenStudy (anonymous):

yes 8 to the power of negative 3 over 10 8 to the power of negative 41 over 30 8 to the power of negative 29 over 30 8 to the power of 71 over 30

OpenStudy (anonymous):

Because according to the problem, I think the answer would be \[\sqrt[5]{8}/32 + \sqrt{2}/16\]

OpenStudy (anonymous):

yikes, I think i'll have to simplify it more

OpenStudy (anonymous):

just a minute

OpenStudy (anonymous):

yeah probably

OpenStudy (anonymous):

ok, so the answer is c

OpenStudy (anonymous):

Are you absolutely positive

OpenStudy (anonymous):

This is because we first have to convert everything to rational exponents and solve. This would look like